OpenStudy (anonymous):

v = s^2 + 1/2sh, solve for h

hero (hero):

\[v = s^2 + \frac{sh}{2}\] 1. subtract \(s^2\) from both sides: \[v - s^2 = \frac{sh}{2}\] 2. multiply both sides by 2: \[2(v - s^2) = sh\] 3. And to finish, divide both sides by s